what is the relation between h andu of a process? what is the value of h-u in the following process?
 
  1. 2SO2(g) +O2(g) -----> 2SO3(g)
  2. H2O(l) -----> H2O(g)
  3. H2(g) + l2(g) -----. 2Hi (g)

Grade:12

1 Answers

Avni Chauhan
5 years ago
\Delta H = \Delta U + \Delta nRT
where

\Delta ndenotes change in the number of gas moles.
i.e. no. of gaseous moles in product - reactant

1)\Delta n= 2-3 = -1
2)\Delta n= 1
3)\Delta n= 2-2 = 0

You can calculate the rest.
P.S. take R = 8.313 J/K mol